# GRADE: Generating multi-hop QA and fine-gRAined Difficulty matrix for RAG Evaluation

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) systems are widely adopted in knowledge-intensive NLP tasks, but current evaluations often overlook the structural complexity and multi-step reasoning required in real-world scenarios. These benchmarks overlook key factors such as the interaction between retrieval difficulty and reasoning depth. To address this gap, we propose GRADE, a novel evaluation framework that models task difficulty along two orthogonal dimensions: (1) reasoning depth, defined by the number of inference steps (hops), and (2) semantic distance between the query and its supporting evidence. We construct a synthetic multi-hop QA dataset from factual news articles by extracting knowledge graphs and augmenting them through semantic clustering to recover missing links, allowing us to generate diverse and difficulty-controlled queries. Central to our framework is a 2D difficulty matrix that combines generator-side and retriever-side difficulty. Experiments across multiple domains and models show that error rates strongly correlate with our difficulty measures, validating their diagnostic utility. GRADE enables fine-grained analysis of RAG performance and provides a scalable foundation for evaluating and improving multi-hop reasoning in real-world applications.
@@
Jeongsoo Lee1*  Daeyong Kwon1,2*  Kyohoon Jin1thanks:   Corresponding author. Email: kyohoon.jin@selectstar.ai 1DATUMO  2Graduate School of Culture Technology, KAIST {jeongsoo.lee, kyohoon.jin}@selectstar.ai, ejmj63@kaist.ac.kr Equal contribution
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has become a widely adopted strategy for enhancing large language models (LLMs) with external knowledge sources Li et al. (2024); Su et al. (2024). By retrieving relevant documents and conditioning generation on this evidence, RAG systems aim to reduce hallucinations and improve factual accuracy in tasks such as open-domain QA, dialogue, and summarization Gao et al. (2023). As LLMs continue to improve, the quality of generation in RAG systems has also advanced. However, the evaluation frameworks used to measure these systems have not evolved at the same pace. Most benchmarks still rely on simple factoid questions that fail to reflect the reasoning complexity and retrieval difficulty faced in real-world scenarios Krishna et al. (2024); Gabburo et al. (2024). Consequently, these datasets offer limited utility in understanding performance variation across tasks.
@@
Current evaluation often lacks the granularity needed to assess the difficulty of individual queries. Tasks are typically treated as uniformly difficult, without considering structural aspects such as the number of reasoning steps involved (e.g., multi-hop reasoning) or the distribution of evidence across multiple documents (e.g., multi-chunk retrieval). Moreover, many existing benchmarks rely on single-hop, single-chunk queries, and define task difficulty using retriever-centric indicators such as lexical ambiguity or document length Gabburo et al. (2024); Salemi and Zamani (2024). Furthermore, collapsing the retrieval and reasoning onto a single difficulty axis complicates the attribution of errors to the retriever, the generator, or their interaction Lee et al. (2025). These coarse-grained approaches hinder nuanced analysis of RAG systems, making it difficult to isolate module-specific failure points—such as retrieval errors versus generation issues Barnett et al. (2024).
@@
Recent work has begun to recognize the importance of task difficulty in evaluating Retrieval-Augmented Generation (RAG) systems. Some studies have proposed taxonomies that categorize queries based on retriever-side metrics such as retrieval recall, query ambiguity, or corpus complexity De Lima et al. (2024). These approaches acknowledge that retrieval difficulty varies across queries and have led to more nuanced analyses of retriever performance. However, such efforts remain largely focused on retrieval alone, without fully considering the downstream reasoning challenges faced by RAG models. Recent work Krishna et al. (2024) has highlighted the importance of evaluating RAG systems in a unified manner that accounts for both retrieval and reasoning. By comparing naïve and multi-step RAG systems using their proposed dataset, the authors demonstrate how evaluation design can directly influence our understanding of a model’s reasoning capabilities.
@@
In particular, most prior work overlooks the structural complexity present in real-world RAG scenarios. Many queries in practice require multi-hop reasoning across multiple semantically diverse sources, which poses significant challenges not only for retrieval but also for evidence integration. In practice, answering a complex question often requires synthesizing information scattered across multiple, semantically distant documents (Lu et al., 2019; De Cao et al., 2019). The cognitive and computational cost of such synthesis grows not just with the number of reasoning steps but also with the semantic dispersion of the supporting evidence. In particular, reasoning across documents from different topical clusters is typically more demanding than connecting closely related passages. For example, answering a multi-hop question like “What legal implications has the use of facial recognition technology had in European countries?” requires synthesizing technical documents on facial recognition systems with legal texts or policy reports from EU jurisdictions.
@@
  • Our framework facilitates more detailed and interpretable analysis of RAG system performance by jointly considering retrieval and generation challenges, thereby supporting more effective error diagnosis and guiding future system improvements.
We release our code and data at https://github.com/anonymous/anonymous to facilitate reproducibility and further research.
@@ 2.3 Multi-Hop QA Evaluation and Fine-Grained Diagnostics
Multi-hop QA tasks require systems to synthesize information scattered across multiple documents or passages. Compared to single-hop queries, they test a model’s ability to perform compositional reasoning, logical inference, and multi-step aggregation. To benchmark such capabilities, a series of datasets have been proposed, each aiming to increase the complexity and diagnostic power of QA evaluation.
@@ 3.1 Knowledge Graph Construction with Missing Link Detection
To ensure that models under evaluation do not have prior exposure to the data, we collected news articles, strictly from July 2024 onward. Articles were split into sentences, and a pretrained fact/opinion classifier was used to retain only factual ones.
To enable reasoning over isolated facts, we applied a decontextualization step: each factual sentence was rewritten into a standalone claim whose meaning is preserved outside the original article. This was done via LLM prompting, followed by a verification stage to ensure semantic consistency with the source sentence. The knowledge graph is composed of entities and relationships extracted from factual claims. Each entity is represented as a node, and each relationship connecting two entities is represented as a directed, labeled edge. A triplet $(e_{i},r,e_{j})$ corresponds to a $(\textit{subject},\textit{predicate},\textit{object})$ structure, mapping directly to a factual claim in the form of a directed relation between entities. We constructed the knowledge graph by merging entities with exact matches across triples.
@@ 3.2 Question Generation
We first enumerated all shortest directed acyclic paths of length 2 to 5 within the knowledge graph and discarded redundant paths that shared the same start and end entities while differing only in intermediate nodes (e.g., $e_{1}\rightarrow e_{3}\rightarrow e_{2}$ vs. $e_{1}\rightarrow e_{4}\rightarrow e_{2}$). From the remaining paths, we randomly sampled 400 per hop count to balance reasoning depth. Afterward, an additional validation step was performed to remove ambiguous QA pairs, resulting in approximately 200 datas per hop. Let $P=\{(e_{1},r_{1},e_{2}),\ldots,(e_{k},r_{k},e_{k+1})\}$ denote a reasoning path of length $k$. We define a Logical Chaining Question (LCQ) $q_{P}$ such that its unique answer is the terminal node $e_{k+1}$, and all supporting facts are contained in the edge set $P$.
@@ 4.1 Experimental Setup
Following prior work, we simulate a realistic Retrieval-Augmented Generation (RAG) scenario in which the knowledge source diverges from the pretraining data of large language models Tang and Yang (2024). To this end, we construct a news dataset using the mediastack API, a RESTful interface that aggregates global news content  [^1] https://mediastack.com/. The dataset comprises articles from a diverse set of English-language sources spanning domains such as sports, health, and science. We focus on articles published between July 2024 and April 2025—a period beyond the knowledge cutoff dates of the LLMs we used—thereby minimizing the likelihood that the base models have seen this content during training. Articles with fewer than 512 tokens were excluded to ensure a minimum level of contextual complexity, while those exceeding 8,192 tokens were removed to avoid excessively large inputs that could hinder processing efficiency. For the RAG database, we split each article into chunks with a total length of 128 to 256 tokens, using a 50-token overlap between consecutive chunks, and embedded them using the OpenAI text-embedding-3-small model.
To evaluate performance across a spectrum of capabilities, we select three representative LLMs: GPT-4o, GPT-4o mini, and o1-mini  [^2] https://openai.com/. GPT-4o is OpenAI’s flagship multimodal model known for its strong performance on a wide range of academic and professional benchmarks. GPT-4o mini, a lightweight variant of GPT-4o, offers high performance with improved speed and cost-efficiency. Lastly, o1-mini is optimized for tasks requiring strong reasoning abilities, such as mathematics and programming. This selection enables a balanced comparison across general-purpose, lightweight, and reasoning-focused LLMs. All prompts used in our experiments are provided in the appendix for transparency and reproducibility.
@@
Table 1: Pearson correlation between retrieval difficulty $D_{r}$ and answer accuracy, computed separately for each fixed hop count, across domains and models.
@@ 4.2 Main Results
To validate the effectiveness of our proposed difficulty framework, we evaluated generator-side, retriever-side, and combined difficulty matrix across three models (o1-mini, GPT-4o, and GPT-4o-mini). We first assessed generator-side difficulty by examining whether increasing the number of hops corresponds to more challenging multi-hop questions. Specifically, we compared the average accuracy across hop counts ranging from 2 to 5. As shown in Figure 2, both the Sports and Health domains exhibited a consistent decrease in average accuracy as the number of hops increased. For example, in the Sports domain, accuracy dropped from 0.68 at 2-hop to 0.48 at 5-hop—a 20% decrease. In the Health domain, accuracy declined from 0.58 to 0.49 over the same hop range, representing a 9% drop. This supports our hypothesis that generator-side difficulty increases with hop count, as answering higher-hop questions typically requires logically connecting more pieces of supporting evidence.
Figure 3: Model performance across varying levels of reasoning depth (2–5 hops) and retrieval difficulty (low vs. high) in both Health and Sports domains. Each heatmap cell represents the error rate for a given setting.
To evaluate retriever-side difficulty, we controlled for generator-side factors by fixing the hop count, ensuring that variations in performance were attributable solely to the retriever. For each fixed hop level, we split the data into four bins based on the retrieval difficulty score $D_{r}$, as defined earlier, and computed the average accuracy for each bin. We then calculated the Pearson correlation between retrieval difficulty and accuracy for each hop level. As shown in Table 1, all models exhibited a strong positive correlation, averaging above 0.6 in the Sports domain and above 0.8 in the Health domain. This indicates that the minimum similarity between the query and its most weakly aligned supporting chunk increases retrieval difficulty $D_{r}$, thereby leading to reduced accuracy.
Finally, to evaluate the combined difficulty framework GRADE that considers both generator-side and retriever-side difficulty, we constructed a $4\times 4$ difficulty matrix where the vertical axis corresponds to reasoning depth (2–5 hops) and the horizontal axis corresponds to retrieval difficulty (divided into four bins based on $D_{r}$ scores). For each cell in the matrix, we computed the average error rate (1 – accuracy) and visualized the results in Figure 3.
@@
Across all models and in both the Sports and Health domains, we observed a general trend: error rates tended to be lower in the upper-left region of the matrix (2-hop, low $D_{r}$) and higher in the bottom-right region (5-hop, high $D_{r}$), with performance degrading steadily along both dimensions in most cases. Notably, values along the diagonal from upper-left to bottom-right showed the steepest increase in error rate, indicating that tasks involving both deeper reasoning and harder retrieval are significantly more difficult. All six matrices exhibited strong linearity across the four diagonal points, with Pearson correlation coefficients exceeding 0.9—highlighting the consistency of this trend. For example, in the Sports domain using the o1-mini model, the error rate increased from 0.241 in the upper-left cell (2-hop, low $D_{r}$) to 0.571 in the bottom-right cell (5-hop, high $D_{r}$), marking a 33% increase. In the Health domain, the difference was even more pronounced, with error rates rising from 0.224 to 0.636—a 41.2% increase. While the overall error rates were higher in Health than in Sports, the gap between easy and hard questions was more prominent in the Sports domain. These results support the validity of our difficulty framework, demonstrating that jointly modeling both generator-side and retriever-side challenges yields an accurate, fine-grained, and diagnostic approach to evaluating multi-hop QA performance in RAG systems.
@@ 4.3 Analysis on Missing Links
In traditional triplet-based multi-hop QA generation pipelines, triplets are connected using shared entities that require an exact match, which often leads to missing links. For example, different expressions referring to the same entity, such as "USA" and "United States," need to be identified and linked. Similarly, cases where phrases like "the Biden administration" and "the U.S. government" refer to the same entity within a specific context also require connection. These missing links represent semantically similar but not explicitly connected entities. To address this, we developed a method described in Section 3.1 that uses semantic clustering and LLMs to supplement the knowledge graph with two types of missing links: exact and contextually equivalent entities. The proportion of data augmented with these missing links, broken down by hop count and domain, is presented in Table 4.

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) systems are widely adopted in knowledge-intensive NLP tasks, but current evaluations often overlook the structural complexity and multi-step reasoning required in real-world scenarios. These benchmarks overlook key factors such as the interaction between retrieval difficulty and reasoning depth. To address this gap, we propose GRADE, a novel evaluation framework that models task difficulty along two orthogonal dimensions: (1) reasoning depth, defined by the number of inference steps (hops), and (2) semantic distance between the query and its supporting evidence. We construct a synthetic multi-hop QA dataset from factual news articles by extracting knowledge graphs and augmenting them through semantic clustering to recover missing links, allowing us to generate diverse and difficulty-controlled queries. Central to our framework is a 2D difficulty matrix that combines generator-side and retriever-side difficulty. Extensive experiments show that error rates strongly correlate with our difficulty measures, validating their diagnostic utility. GRADE enables fine-grained analysis of RAG performance and provides a scalable foundation for evaluating and improving multi-hop reasoning in real-world applications.
@@
Jeongsoo Lee1*   Daeyong Kwon1,2*   Kyohoon Jin1thanks: Corresponding author. 1DATUMO  2Graduate School of Culture Technology, KAIST {jeongsoo.lee, kyohoon.jin}@selectstar.ai, ejmj63@kaist.ac.kr Equal contribution
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has become a widely adopted strategy for enhancing large language models (LLMs) with external knowledge sources Li et al. (2024); Su et al. (2024). By retrieving relevant documents and conditioning generation on this evidence, RAG systems aim to reduce hallucinations and improve factual accuracy in tasks such as open-domain QA, dialogue, and summarization Gao et al. (2023). As LLMs continue to improve, the quality of generation in RAG systems has also advanced. However, the evaluation frameworks used to measure these systems have not kept pace. Most benchmarks still rely on simple factoid questions that fail to capture the reasoning complexity and retrieval difficulty faced in real-world scenarios Krishna et al. (2024); Gabburo et al. (2024). Consequently, these datasets offer limited utility in understanding performance variation across tasks.
@@
Current evaluation often lacks the granularity needed to assess query difficulty. Tasks are typically treated as uniformly difficult, without considering structural aspects such as the number of reasoning steps (multi-hop reasoning) or the distribution of evidence across multiple documents (multi-chunk retrieval). Moreover, many benchmarks rely on single-hop, single-chunk queries, and define difficulty using retriever-centric indicators such as lexical ambiguity or document length Gabburo et al. (2024); Salemi and Zamani (2024). Furthermore, collapsing retrieval and reasoning onto a single axis complicates the attribution of errors to the retriever, the generator, or their interaction Lee et al. (2025). These coarse-grained approaches hinder nuanced analysis of RAG systems, making it difficult to isolate module-specific failure points—such as retrieval errors versus generation issues Barnett et al. (2024).
@@
Recent work has begun to recognize the importance of task difficulty in evaluating RAG systems. Some studies have proposed taxonomies that categorize queries based on retriever-side metrics such as retrieval recall, query ambiguity, or corpus complexity De Lima et al. (2024). These approaches acknowledge that retrieval difficulty varies across queries and have led to more nuanced analyses of retriever performance. However, such efforts remain largely focused on retrieval alone, without fully considering the downstream reasoning challenges faced by RAG models. Recent work Krishna et al. (2024) has highlighted the importance of evaluating RAG systems in a unified manner that accounts for both retrieval and reasoning. By comparing naïve and multi-step RAG systems using their proposed dataset, the authors demonstrate how evaluation design can directly influence our understanding of a model’s reasoning capabilities.
@@
In particular, most prior work overlooks the structural complexity in real-world RAG scenarios. Many queries in practice require multi-hop reasoning across multiple semantically diverse sources, which poses significant challenges for retrieval and evidence integration. In practice, answering a complex question often requires synthesizing information scattered across multiple, semantically distant documents (Lu et al., 2019; De Cao et al., 2019). The cognitive and computational cost of such synthesis grows not just with the number of reasoning steps but also with the semantic dispersion of the supporting evidence. In particular, reasoning across documents from different topical clusters is typically more demanding than connecting closely related passages. For example, answering a multi-hop question like “What legal implications has the use of facial recognition technology had in European countries?” requires synthesizing technical documents on facial recognition systems with legal texts or policy reports from EU jurisdictions.
@@
  • Our framework facilitates more detailed and interpretable analysis of RAG system performance by jointly considering retrieval and generation challenges, thereby supporting more effective error diagnosis and guiding future system improvements, as illustrated in Figure 1.
We release our code to facilitate reproducibility and further research.  [^1] https://github.com/DaeyongKwon98/GRADE
@@ 2.3 Multi-Hop QA Evaluation and Fine-Grained Diagnostics
Multi-hop QA tasks require systems to synthesize information scattered across multiple documents. Compared to single-hop queries, they test a model’s ability to perform compositional reasoning, logical inference, and multi-step aggregation. To benchmark such capabilities, a series of datasets have been proposed, each aiming to increase the complexity and diagnostic power of QA evaluation.
@@ 3.1 Knowledge Graph Construction with Missing Link Detection
To ensure that models under evaluation do not have prior exposure to the data, we collected news articles, strictly from July 2024 onward. Articles were split into sentences, and a pretrained fact/opinion classifier was used to retain only factual ones. To enable reasoning over isolated facts, we applied a decontextualization step: each sentence was rewritten into a standalone claim whose meaning is preserved outside the original article. This was done via prompting, followed by a verification stage to ensure semantic consistency with the source sentence. The knowledge graph is composed of entities and relationships extracted from factual claims. Each entity is represented as a node, and each relationship connecting two entities is represented as a directed, labeled edge. A triplet $(e_{i},r,e_{j})$ corresponds to a $(\textit{subject},\textit{predicate},\textit{object})$ structure, mapping directly to a factual claim in the form of a directed relation between entities. We constructed the knowledge graph by merging entities with exact matches across triples.
@@ 3.2 Question Generation
We first enumerated all shortest directed acyclic paths of length 2 to 5 within the knowledge graph and discarded redundant paths that shared the same start and end entities while differing only in intermediate nodes. From the remaining paths, we sampled 400 per hop count to balance reasoning depth. Afterward, an additional validation step was done to remove ambiguous QA pairs, resulting in approximately 200 datas per hop. Let $P=\{(e_{1},r_{1},e_{2}),\ldots,(e_{k},r_{k},e_{k+1})\}$ denote a reasoning path of length $k$. We define a Logical Chaining Question (LCQ) $q_{P}$ as a question for which all supporting facts are contained within the edge set $P$.
@@ 3.4 RAG Evaluation
While this setup corresponds to a standard single-step RAG, we also conducted additional experiments, reported in Appendix A.4, with Multi-step RAG and Query-decomposition RAG. These variants address more complex information needs by enabling iterative retrieval and refinement or by decomposing questions into simpler sub-queries. Incorporating them allows us to evaluate whether more advanced retrieval strategies yield improvements beyond the baseline.
@@ 4.1 Experimental Setup
Following prior work, we simulate a realistic RAG scenario in which the knowledge source diverges from the pretraining data of large language models Tang and Yang (2024). To this end, we construct a news dataset using the mediastack API, a RESTful interface that aggregates global news content  [^2] https://mediastack.com/. The dataset comprises articles from a diverse set of English-language sources spanning domains such as sports, health, and science. We focus on articles published between July 2024 and April 2025—a period beyond the knowledge cutoff dates of the LLMs we used—thereby minimizing the likelihood that the base models have seen this content during training. Articles with fewer than 512 tokens were excluded to ensure a minimum level of contextual complexity, while those exceeding 8,192 tokens were removed to avoid excessively large inputs that could hinder processing efficiency. For the RAG database, we split each article into chunks with a total length of 128 to 256 tokens, using a 50-token overlap between consecutive chunks. We embedded them using four representative embedding models: OpenAI text-embedding-3-small, all-mpnet-base-v2, bge-en-large-v1.5, and jina-embeddings-v3. This selection covers both closed-source and open-source approaches, ensuring a balanced and comprehensive evaluation across different embedding paradigms.
To evaluate performance across a spectrum of capabilities, we select five representative LLMs: GPT-4o, GPT-4o mini, o1-mini, Claude-4-Sonnet, and Llama 3.2 3B Instruct  [^3] https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/. This selection enables a balanced comparison across general-purpose, lightweight, reasoning-focused, and open-source LLMs. All prompts are provided in the appendix A.9 for transparency and reproducibility.
@@
Table 1: Pearson correlation between retrieval difficulty $D_{r}$ and error rate, computed separately for each fixed hop count, across domains and models.
@@ 4.2 Main Results
To validate the effectiveness of our proposed difficulty framework, we evaluated generator-side, retriever-side, and combined difficulty matrix. We first assessed generator-side difficulty by examining whether increasing the number of hops corresponds to more challenging multi-hop questions. Specifically, we compared the average accuracy across hop counts ranging from 2 to 5. As shown in Figure 2, both the Sports and Health domains exhibited a consistent decrease in average accuracy as the number of hops increased. For example, in the Sports domain, accuracy dropped from 0.68 at 2-hop to 0.48 at 5-hop—a 20% decrease. In the Health domain, accuracy declined from 0.58 to 0.49 over the same hop range, showing a 9% drop. This supports our hypothesis that generator-side difficulty increases with hop count, as answering higher-hop questions typically requires logically connecting more pieces of supporting evidence.
Figure 3: Error rates (1-accuracy) across varying levels of reasoning depth (2–5 hops) and retrieval difficulty (low vs. high) in both Health and Sports domains.
To evaluate retriever-side difficulty, we controlled for generator-side factors by fixing the hop count, ensuring that variations in performance were attributable solely to the retriever. For each fixed hop level, we split the data into four bins based on the retrieval difficulty score $D_{r}$, as defined earlier, and computed the average accuracy for each bin. We then calculated the Pearson correlation between retrieval difficulty and error rate for each hop level. As shown in Table 1, all models exhibited a strong positive correlation, averaging above 0.6 in the Sports domain and above 0.8 in the Health domain. This indicates that the minimum similarity between the query and its most weakly aligned supporting chunk increases retrieval difficulty $D_{r}$, thereby leading to reduced accuracy.
Finally, to evaluate the combined difficulty framework GRADE that considers both generator-side and retriever-side difficulty, we constructed a $4\times 4$ difficulty matrix where the vertical axis corresponds to reasoning depth (2–5 hops) and the horizontal axis corresponds to retrieval difficulty (divided into four bins based on $D_{r}$ scores). For each cell in the matrix, we computed the average error rate and visualized the results in Figure 3.
@@
Across all models and in both the Sports and Health domains, we observed a general trend: error rates tended to be lower in the upper-left region of the matrix (2-hop, low $D_{r}$) and higher in the bottom-right region (5-hop, high $D_{r}$), with performance degrading steadily along both dimensions in most cases. Notably, values along the diagonal from upper-left to bottom-right showed the steepest increase in error rate, indicating that tasks involving both deeper reasoning and harder retrieval are significantly more difficult. All matrices exhibited strong linearity across the four diagonal points, with Pearson correlation coefficients exceeding 0.9—highlighting the consistency of this trend. For example, in the Sports domain with o1-mini, the error rate increased from 0.241 in the upper-left cell (2-hop, low $D_{r}$) to 0.571 in the bottom-right cell (5-hop, high $D_{r}$), marking a 33% increase. In the Health domain, the difference was even more pronounced, with error rates rising from 0.224 to 0.636—a 41.2% increase. While the overall error rates were higher in Health than in Sports, the gap between easy and hard questions was more prominent in the Sports domain. These results support the validity of our difficulty framework, demonstrating that jointly modeling both generator-side and retriever-side challenges yields an accurate, fine-grained, and diagnostic approach to evaluating multi-hop QA performance in RAG systems. The results of our experiments with additional models, embeddings, RAG systems, domains, and chunk sizes are reported in the Appendix A.
@@ 4.3 Analysis on Missing Links
In traditional triplet-based multi-hop QA generation pipelines, triplets are connected using shared entities that require an exact match, which often leads to missing links. For example, different expressions referring to the same entity, such as "USA" and "United States," need to be identified and linked. Similarly, cases where phrases like "Biden administration" and "U.S. government" refer to the same entity within a specific context also require connection. These missing links represent semantically similar but not explicitly connected entities. To address this, we developed a method described in Section 3.1 that uses semantic clustering and LLMs to supplement the knowledge graph with two types of missing links: exact and contextually equivalent entities. The proportion of data augmented with these missing links, broken down by hop count and domain, is presented in Table 4.